one angle of a triangle has a measure that is 10 more than twice that of the second angle. the third angle has a measure 10 less than the first. find the measure of each angle of the triangle
Answer by checkley71(8403)

X=2Y+10
Z=X-10
X+Y+Z=180 REPLACE X WITH(2Y+10) & Z WITH(2Y+10-10)
2Y+10+Y+2Y+10-10=180
5Y=180-10
5Y=170
Y=170/5
Y=34 ANSWER FOR THE SECOND ANGLE.
X=2*34+10=68+10=78 ANSWER FOR THE FIRST ANGLE.
Z=78-10=68 FOR THE THIRD ANGLE.
PROOF:
34+78+68=180
180=180